OpenStudy (anonymous):

Simplify. Do not use negative exponents in the answer. Y−5 · y

OpenStudy (anonymous):

it should actually read y^-5y

OpenStudy (anonymous):

any number raised to a negative power is equivalent to 1 divided by that number raised to a positive power

OpenStudy (anonymous):

x^-y = 1/x^y

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

sure, would you like to write the answer so i can check it?

OpenStudy (anonymous):

sure give me a minute pls

OpenStudy (anonymous):

y^-5y = y5y= y^6 ?

OpenStudy (anonymous):

not quite

OpenStudy (anonymous):

ohhh pls help me understand more

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

see how i said that x^-y = 1/x^y ?

OpenStudy (anonymous):

alll you need to do is put the number in the denominator and get rid of the negative sign in front of the exponent

OpenStudy (anonymous):

y^y/5 ?

OpenStudy (anonymous):

a litter closer

OpenStudy (anonymous):

5/y^y?

OpenStudy (anonymous):

do I add the 'y's?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

you ignore the exponents for now and concentrate on only the negative sign

OpenStudy (anonymous):

for example, 4^-1 = 1/4^1 = 1/4

OpenStudy (anonymous):

so is it 1/y^5y?

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

ok ok

OpenStudy (anonymous):

you have the right answer

OpenStudy (anonymous):

ohhhh tytytyty :)

OpenStudy (anonymous):

just for another example for when you see fractions with negative exponents. x^-x/y^-y = y^y/x^x

OpenStudy (anonymous):

just reverse the numerator and denominator and remove the negative sign from the exponent